# Understanding How Psychological Distance Influences User Preferences in Conversational Versus Web Search

### v1

@@ 4.2.2. Unique Themes for Preferring WebSearch
WebSearch has fewer unique themes than ConvSearch, but nevertheless has advantages in certain areas.
@@
Specifically, for time-sensitive information, WebSearch is uniquely able to help users verify timeliness. As one participant commented: “I’d rather use the web search so I… can see how recent the information is.” (P121)
@@ 4.2.3. Themes Found in Both Systems
Others (n = 12) expressed contrary views, believing that web search results might contain more misinformation, making conversational search more reliable and accurate. One noted: “I would choose the chatbot rather than a web search because I would not have to look through multiple web pages for information and potentially run into misinformation and instead I can get trustworthy answers from the chatbot itself.” (P117)
Discovery and Exploration. WebSearch’s result list offered participants (n = 4) greater freedom to explore information and gain new ideas and interests during the exploration process. One participant said: “I would rather use the web search because I like to look things up and decide which sites I want to use for information on my own. Also, when you do a web search you may find things that spark new ideas and interests. I think it might be useful to use a chatbot to get quick bits of information, and the use the web to search for the things you are most interested in.” (P18)
For some participants (n = 9), the more comprehensive information provided by ConvSearch served as a foundation to gain new ideas and further explore the information. One participant said: “I think it’s a good foundational point to work off of, especially when creating an itinerary at a location I know nothing about. Chatbot gave me ideas of what to explore further and potentially search to see reviews before finalizing.” (P116)
Combining Two Tools  [^7] Detailed data can be found in the Appendix 7.3.. Some participants (n = 9), despite being explicitly instructed to provide reasons for their preference for either web search or conversational search, still suggested combining both tools. As one noted: “Chatbot makes it easier to get a lot of info as a great starting point. I would then use a search engine to verify the info from the chatbot.” (P32)
Comprehensive Information. For planning trips, one participant (n = 1) preferred to refer to real experiences shared by other travelers. WebSearch can provide these authentic experiences, which ConvSearch cannot. Therefore, WebSearch offers more comprehensive information. “I would prefer a search box because it gives more details and articles from previous experiences.” (P43)
@@ 5. Discussion
Our findings highlight the pivotal role of psychological distance of information seeking tasks in shaping user perceptions and preferences toward different IR technologies. As psychological distance increased, participants showed a preference for LLM-powered conversational search over web search, perceiving it as easier to use, more credible, useful, and enjoyable. Consequently, user acceptance of LLM-powered conversational search also rose with increasing psychological distance. These critical insights highlight how the ubiquitous factor of psychological distance informs design considerations for the nascent technology that is generative AI-based conversational search.
@@ 5.1. Deep Impact of Psychological Distance in Information Seeking Tasks (RQ)
Our findings corroborate CLT’s suggested existence of a relationship between psychological distance and user evaluations, and hence choice of IR technologies (Liberman and Trope, 2014; Trope and Liberman, 2010; Trope et al., 2007). People’s choices and preferences are based on their understanding of events, which is in turn affected by psychological distance (Trope and Liberman, 2010). Our research provides a comprehensive view from the nature of tasks to user preferences, not only enhancing our understanding of the effects of task psychological distance but also offering practical guidance for future IR system design. By addressing user information needs at different psychological distances, IR systems can further improve user interaction experiences and satisfaction.
@@ 5.1.1. Greater Psychological Distance Makes Conversational Search Easier to Use (H1)
Our results suggest that tasks with greater psychological distance make participants find conversational search easier than web search. This is due to conversational search providing integrated and summarized information tailored to specific questions (Kaushik and Jones, 2023). It aligns better with the high-level abstract thinking used by users in these situations (Hadar et al., 2022; Yoo and Lee, 2023). Conversely, web search provides a list of results, requiring users to filter and summarize information themselves.
Our thematic analysis supports this perspective. Participants cited themes Low Workload and Specific Information Consolidation for their preference for conversational search.Low Workload is linked to its perceived ease of use, as less workload allows users to meet their needs more easily. Specific Information Consolidation is seen as a reason for Low Workload, illustrating how conversational search’s ability to integrate information underpins its ease of use.
@@ 5.1.2. Greater Psychological Distance Makes Conversational Search More Credible (H2)
Further analysis of the three dimensions of credibility  [^8] Specific data are presented in Appendix 7.4. reveals that although user perceptions of the accuracy and believability of information from LLM-powered conversational search significantly improve with greater psychological distance, the perceived enhancement in information authenticity is not as pronounced. This may be because assessing information authenticity depends more on directly verifying information sources, an area where LLM-powered conversational search is weaker compared to web search (Bender et al., 2021). Additionally, conversational search is limited by its LLM core, which is prone to hallucinations (Maynez et al., 2020) and relies on outdated training data (Achiam et al., 2023), further reducing user perception of the authenticity of the information provided. Therefore, LLM-powered conversational search still needs to strengthen its capabilities regarding verifiable information sources, hallucination issues, and providing up-to-date information to address concerns about information authenticity.
5.1.3. Greater Psychological Distance Makes Conversational Search More Useful and Enjoyable (H3-H4)
As psychological distance is far, user perceptions of the usefulness and enjoyment of LLM-powered conversational search improve significantly. This underscores that increasing the psychological distance of a task can enhance user perceptions of the usefulness and enjoyment of conversational search, even without changes in the technology itself. This finding aligns with the TAM and its extensions (Davis, 1989; Davis et al., 1992, 1992), which shows the positive interlink between perceived ease of use (Adams et al., 1992), credibility (Ayeh, 2015), enjoyment, and usefulness. In our thematic analysis, Speed and Efficiency are directly related to user perception of usefulness. Participants find conversational search useful because it allows them to complete tasks faster and more efficiently than web search. The perception of enjoyment of conversational search results from enhanced interactivity through natural dialogue (theme Interactive and Conversational). Some participants feel that interacting with conversational systems is like conversing with another human, aligning with research findings that humanlike qualities are a significant factor in making chatbots enjoyable (Følstad and Brandtzaeg, 2020).
@@ 5.2. Overall User Acceptance of Conversational Search (H5)
Our findings align with the TAM and its extensions (Davis, 1989; Davis et al., 1992), suggesting that improvements in the previous four user metrics culminate in increased user acceptance of LLM-powered IR systems as psychological distances increase (H5). Additionally, we found that controlling for psychological distance, users generally prefer LLM-powered conversational search over traditional web search. For common information seeking tasks like travel planning, this preference may stem from the inherent advantages of LLM-powered conversational search in areas such as Low Workload, Speed and Efficiency, Specific Information Consolidation, and Interactive and Conversational capabilities. These characteristics enable LLM-powered conversational search systems to more effectively meet users’ information needs in daily tasks.
@@ 5.3. Design Implications
Our research confirms the significant impact of the psychological distance of information seeking tasks on user perception and preferences. Conversational search, with its ability to provide information from an abstract perspective, better meets user needs when the psychological distance is greater. However, current conversational search systems exhibit a fixed level of abstraction in information provision, regardless of human cognitive state (Kirshner, 2024). This limits the effectiveness of LLMs in providing information, as it does not fully meet the diverse judgment needs inherent in human cognition (Santurkar et al., 2023). Past HCI research has also touched on similar views when discussing the design of conversational search. They identified user demands for different levels of information, with some preferring direct access to answers, while others favored receiving broad background information (Vtyurina et al., 2017). This aligns with the two information levels that we have elucidated. However, these studies did not identify the key factors behind these differing information needs.
We propose using psychological distance as a primary guiding factor to enhance the diversity of information levels provided by conversational search systems. By considering psychological distance, we can not only increase the diversity of information, but also better meet the information needs at the user’s current cognitive state.
@@ 5.3.1. Impact on Conversational Search Design
In further enhancing LLM-powered conversational search systems, it is crucial to fully leverage the capabilities of LLMs by assessing the psychological distance of information seeking tasks through dialogue content and treating it as a key attribute. For example, the system could evaluate the psychological distance of tasks by analyzing user descriptions across the spatial, temporal, social, and hypothetical dimensions. Additionally, this assessment could use the capability of conversational search to pose supplementary questions (Radlinski and Craswell, 2017), refining the evaluation of task psychological distance through additional inquiries focused on these dimensions.
Conversational search systems could adjust their IR strategies based on the psychological distance perceived by users. For tasks with a closer psychological distance, users typically seek specific and actionable information. In such cases, search results should focus on providing detailed steps and practical guidance. For example, when planning a trip in the near future, the system could provide specific travel itineraries, recommended accommodations, and dining options. This approach not only satisfies the user’s low-level, specific information needs but also aids in making quick decisions and executing plans.
@@ 5.3.2. Combine Conversational Search and Web Search
Web search, a widely used IR technology, offers significant advantages in visualization and source reliability that should not be overlooked. Currently, clear guidance on integrating conversational search with web search is lacking, often leaving users to choose their preferred tool without knowing which would better satisfy their information needs. We suggest offering tools that more closely align with current information needs based on the psychological distance of a given task. Therefore, we propose several recommendations for systems that integrate conversational and web search.
@@
For tasks with closer psychological distances, we recommend returning a list of web search results. This is because when the task’s psychological distance is close, users typically need specific and detailed information, and web search can provide a rich and detailed list of sources. For example, when searching for local restaurants or immediate weather forecasts, web search can provide timely and specific information that directly meets users’ needs.
@@
However, for tasks with a greater psychological distance, we suggest providing synthesized information generated by conversational search. In these cases, users tend to seek overview or conceptual information, and conversational search can meet these needs by integrating and summarizing information. For example, when exploring potential travel destinations or learning about new technologies, conversational search can provide a comprehensive information framework to help users build a preliminary understanding of the topic.
@@
Although past HCI research has not focused on the psychological distance of tasks, it has explored factors influencing user choices between conversational and web search in specific task scenarios. A study on programming tasks noted that programmers opt for conversational search when objectives are clear, seeking results aligned with their goals (Yen et al., 2024). This phenomenon, stemming from the confirmatory nature of conversational search, known as the echo chamber effect, has been extensively studied in HCI (Sharma et al., 2024; Tanprasert et al., 2024). This characteristic of conversational search limits users’ exposure to diverse viewpoints and is difficult to fundamentally resolve due to its LLM core. To optimize this, some studies have suggested designs that allow conversational systems to actively present different viewpoints (Feltwell et al., 2020) and designs that proactively remind users of current information biases in conversations (Sharma et al., 2024). However, our proposed solution is to integrate conversational and web search, leveraging the diversity of viewpoints offered by web search to balance the confirmatory bias of conversational search. Additionally, HCI research discussing user trust in ChatGPT versus web search noted that generative AI’s lack of information sources reduces user trust (Jung et al., 2024). Yet, its assistance in ideation encourages its use. Our participants expressed similar views [^9] Specific data are presented in Appendix 7.3., preferring to obtain preliminary information through conversational search, then verify this information and acquire more details through web search.
Therefore, we propose a deeper integration of the two tools. Even when providing synthesized information from conversational search, users should be allowed to click on keywords or content mentioned in the synthesis to directly transition to web search. Through this approach, our goal is to use the web search to help users verify the accuracy of related information, while retaining the benefits of conversational search.
@@ 5.4. Limitations
This study has some limitations. First, our study only considered information seeking tasks in travel scenarios. These tasks, being general in nature, may have influenced user perceptions of LLM-powered conversational search, which is better suited for general rather than specialized tasks. This limitation might obscure the specific differences in user perceptions between web search and LLM-powered conversational search. For instance, we observed no significant differences in most metrics when the psychological distance was near. This pattern may not apply to other scenarios and also prevented us from further exploring the specific impacts of tasks with near psychological distance on both types of IR tools. However, with web search as a control, we found no significant differences in most ratings between the two IR tools when the psychological distance was near, but significant differences emerged when the psychological distance was far. This outcome still addresses our main research question and reveals a general principle: Users prefer LLM-powered conversational search when the psychological distance is greater.
Furthermore, the custom search interface we implemented does not fully represent a real, commercial search page; it lacks features like image previews and related searches, especially compared to the actual Google Search website. However, since the search results are sourced from the genuine Google Search API, we consider them to still fairly represent real search interactions.

### v2

@@ 5.2.2. Unique Themes for Preferring WebSearch
WebSearch has fewer unique themes than ConvSearch, but it demonstrates the irreplaceable advantages of WebSearch.
@@
Specifically, for time-sensitive information, WebSearch is uniquely able to help users verify timeliness. As one participant commented: “I’d rather use the web search so I can see the sources of where I’m getting my information from. Also can see how recent the information is.” (P121)
@@ 5.2.3. Themes Found in Both Systems
Some participants (n = 12) expressed contrary views, believing that web search results might contain more misinformation, making conversational search more reliable and accurate. One noted: “I would choose the chatbot rather than a web search because I would not have to look through multiple web pages for information and potentially run into misinformation and instead I can get trustworthy answers from the chatbot itself.” (P117)
Discovery and Exploration. WebSearch’s result list offers participants (n = 4) greater freedom to explore information and gain new ideas and interests during the exploration process. One participant said: “I would rather use the web search because I like to look things up and decide which sites I want to use for information on my own. Also, when you do a web search you may find things that spark new ideas and interests. I think it might be useful to use a chatbot to get quick bits of information, and the use the web to search for the things you are most interested in.” (P18)
For some participants (n = 9), the more comprehensive information provided by ConvSearch serves as a foundation to gain new ideas and further explore the information. One participant said: “I think it’s a good foundational point to work off of, especially when creating an itinerary at a location I know nothing about. Chatbot gave me ideas of what to explore further and potentially search to see reviews before finalizing.” (P116)
Combining Two Tools  [^9] Detailed data can be found in the Appendix (A.3), uploaded with Supplementary Material.. Some participants (n = 9), despite being explicitly instructed to provide reasons for their preference for either web search or conversational search, still suggested combining both tools. As one noted: “Chatbot makes it easier to get a lot of info as a great starting point. I would then use a search engine to verify the info from the chatbot.” (P32)
Comprehensive Information. For planning trips, one participant prefers to refer to real experiences shared by other travelers. WebSearch can provide these authentic experiences, which ConvSearch cannot. Therefore, WebSearch offers more comprehensive information. “I would prefer a search box because it gives more details and articles from previous experiences.” (P43)
@@ 6. Discussion
Our findings highlight the important role of psychological distance of information seeking tasks in shaping user perceptions and preferences toward different IR technologies. As psychological distance increased, participants showed a preference for LLM-powered conversational search over web search, under fixed condition ordering where web search preceded conversational search. They perceived conversational search as easier to use, more credible, useful, and enjoyable. Consequently, user acceptance of LLM-powered conversational search also rose with increasing psychological distance. These valuable insights highlight how the ubiquitous factor of psychological distance informs design considerations for the nascent technology that is generative AI-based conversational search.
@@ 6.1. Deep Impact of Psychological Distance in Information Seeking Tasks (RQ)
Our findings support CLT’s suggested existence of a relationship between psychological distance and user evaluations, and hence choice of IR technologies (Liberman and Trope, 2014; Trope and Liberman, 2010; Trope et al., 2007). People’s choices and preferences are based on their understanding of events, which is in turn affected by psychological distance (Trope and Liberman, 2010). Our research provides a comprehensive view from the nature of tasks to user preferences, not only enhancing our understanding of the effects of task psychological distance but also offering practical guidance for future IR system design. By addressing user information needs at different psychological distances, IR systems can further improve user interaction experiences and satisfaction.
@@ 6.1.1. Greater Psychological Distance Makes Conversational Search Easier to Use (H1)
Our results suggest that with tasks with greater psychological distance, participants find conversational search easier to use than web search. Given that far-distance users tend to think in more abstract and general concepts (Trope et al., 2007), this preference is likely explained by conversational search better aligning with information needs for these participants. LLMs have been shown to be adaptable to the construal level required by a user (Yoo and Lee, 2023), and can provide integrated and summarized information tailored to specific questions (Kaushik and Jones, 2023). Search engines, naturally, lack this capability, and require users to manually filter and summarize information from a list of results. Far-distance tasks that require high-level and abstract information are therefore more easily accomplished with LLM-powered search.
Our thematic analysis also supports this perspective. Participants cited Low Workload and Specific Information Consolidation as reasons for preferring conversational search. The Low Workload theme is directly linked to perceived ease of use, as less workload allows users to meet their needs more easily. Specific Information Consolidation can also be linked to Low Workload in this context, illustrating how conversational search’s ability to integrate information underpins its ease of use.
@@ 6.1.2. Greater Psychological Distance Makes Conversational Search More Credible (H2)
Further analysis of the three dimensions of credibility [^10] Specific data are presented in the Appendix (A.4), uploaded with Supplementary Material. reveals that although user perceptions of the accuracy and believability of conversational search tend to improve with greater psychological distance, the perceived enhancement in information authenticity is less pronounced. Unlike accuracy and believability, authenticity is related more to the legitimacy of the source of a message, rather than the factuality of the message content itself (Kang, 2010). In the context of our travel task, authentic content might come from, for instance, a passionate and well-liked travel blogger. LLMs presently still struggle to reliably provide source attributions (Maynez et al., 2020; Bender et al., 2021), and themselves are unlikely to be regarded as ”authentic” as real human writers; this may account for the fact that they score relatively lower on this metric. We note, however, that this may vary with the nature of a given task: certain tasks, particularly those involving searching for lived human experiences, may emphasize perceived authenticity more than other tasks relying on purely factual information.
Recent advancements in AI-powered search systems, such as OpenAI’s integration of web browsing capabilities (OpenAI, 2024), have begun to address some limitations of standalone conversational or web search tools. These systems now offer source attribution and access to real-time information, significantly enhancing their utility. However, challenges remain in dynamically adapting to different user tasks and mitigating the confirmatory bias inherent in conversational search.
6.1.3. Greater Psychological Distance Makes Conversational Search More Useful and Enjoyable (H3, H4)
As psychological distance increases, user perceptions of the usefulness and enjoyment of LLM-powered conversational search improve significantly. This underscores that increasing the psychological distance of a task can enhance user perceptions of the usefulness and enjoyment of conversational search, even without changes in the technology itself. This finding aligns with the TAM and its extensions (Davis, 1989; Davis et al., 1992), which shows the positive interlink between perceived ease of use (Adams et al., 1992), credibility (Ayeh, 2015), enjoyment, and usefulness. In our thematic analysis, Speed and Efficiency are directly related to user perception of usefulness. Participants find conversational search useful because it allows them to complete tasks faster and more efficiently than web search. The perception of enjoyment of conversational search results from enhanced interactivity through natural dialogue (Interactive and Conversational). Some participants feel that interacting with conversational systems is like conversing with another human, aligning with research findings that humanlike qualities are an influential factor in making chatbots enjoyable (Følstad and Brandtzaeg, 2020).
@@ 6.2. Overall User Acceptance of Conversational Search (H5)
Our findings align with the TAM and its extensions (Davis, 1989; Davis et al., 1992, 1992), suggesting that improvements in the previous four user metrics culminate in increased user acceptance of LLM-powered IR systems as psychological distances increase (H5). Additionally, we found that controlling for psychological distance, users generally prefer LLM-powered conversational search over traditional web search. For common information seeking tasks like travel planning, this preference may stem from the inherent advantages of LLM-powered conversational search in areas such as Low Workload, Speed and Efficiency, Specific Information Consolidation, and Interactive and Conversational capabilities. These characteristics enable LLM-powered conversational search systems to more effectively meet users’ information needs in daily tasks.
Another important aspect influencing this increased acceptance is the perceived credibility of LLM-powered conversational search systems, particularly when psychological distance is high. Unlike other factors, which directly influence user acceptance, credibility exerts its impact on acceptance indirectly. As discussed in Section 6.1.2, tasks with greater psychological distance lead to higher perceived information credibility for conversational search compared to web search. For example, some participants reported that conversational search is more reliable, as it filters results and provides accurate responses. This aligns with prior research showing that perceived credibility enhances user trust in automated systems by alleviating concerns about misinformation or irrelevant results (Fogg and Tseng, 1999; Kim et al., 2008). When users perceive a system as credible, they are more likely to view it as reliable and efficient, fostering greater acceptance. However, our findings also highlight nuances in how credibility influences acceptance. While conversational search excels in perceived accuracy and believability, its perceived authenticity remains lower compared to web search, as LLMs often lack robust source attributions (Maynez et al., 2020; Bender et al., 2021). This limitation may partially constrain user acceptance in tasks where source legitimacy is critical, such as those requiring lived human experiences or expert opinions. Nevertheless, for tasks with greater psychological distance—where users prioritize abstract or general information over specific, source-authenticated details—credibility perceptions of conversational search systems are sufficiently high to drive overall acceptance.
@@ 6.3. Design Implications
Our research has confirmed the impact of psychological distance in information seeking tasks on user perception and preferences. Conversational search, with its ability to provide information from an abstract perspective, better meets user needs when the psychological distance is greater. However, current conversational search systems exhibit a fixed level of abstraction in information provision, regardless of human cognitive state (Kirshner, 2024). This limits the effectiveness of LLMs in providing information, as it does not fully meet the diverse judgment needs inherent in human cognition (Santurkar et al., 2023). Past HCI research has also touched on similar views when discussing the design of conversational search, as they identified user demands for different levels of information, with some preferring direct access to answers, while others favored receiving broad background information (Vtyurina et al., 2017). This aligns with the two information levels that we have elucidated. However, these studies did not identify the primary factors influencing user information needs.
We propose using psychological distance as a consideration to enhance the diversity of information levels provided by conversational search systems. By considering psychological distance, we can not only increase the diversity of information, but also better meet the information needs at the user’s current cognitive state.
@@ 6.3.1. Impact on Conversational Search Design
To leverage the capabilities of LLM-powered conversational search systems, it is important to assess the psychological distance of information seeking tasks through dialogue content and treat it as a significant attribute. An individual’s current psychological distance to a task manifests in the way they describe events or actions. This involves omitting features and details as distance increases, e.g. ”having breakfast in a Parisian café by the Riviera” vs. ”a romantic vacation in France” (Trope et al., 2007). Using LLMs or other natural language processing means, IR systems can automatically analyze a user’s inputs for either concrete details or abstract concepts, to determine the construal level (and hence psychological distance) a user is operating at. Additionally, systems may also use the capability of conversational search (Radlinski and Craswell, 2017) to pose supplementary questions about the user’s goals along the four dimensions of spatial, temporal, social, and hypothetical distance, further refining the evaluation of psychological distance.
Conversational search could adjust its IR strategy based on the psychological distance perceived by users. For tasks with a closer psychological distance, users typically seek specific and actionable information. In such cases, search results should focus on providing detailed steps and practical guidance. For example, when planning a trip, the system could provide specific travel itineraries, recommended accommodations, and dining options. This approach not only satisfies the user’s low-level, specific information needs but also aids in making quick decisions and executing plans.
@@
Additionally, to further optimize the performance of LLM-powered conversational search, we also provide recommendations for fine-tuning LLMs specifically for IR. We suggest labeling internet data used for model fine-tuning to distinguish between specific and abstract information in response to queries. Through such fine-tuning, the model can more accurately understand the level of information, thereby providing more personalized and precise search results. Implementing these recommendations, the optimized conversational search will not only better meet user information needs but also enhance the search experience and strengthen user preferences.
@@ 6.3.2. Combine Conversational Search and Web Search
Currently, clear guidance on integrating conversational search with web search is lacking, often leaving users to choose their preferred tool without knowing which would better satisfy their information needs. Considering the psychological distance of tasks, it is advisable to proactively offer tools that more closely align with current information needs based on this criterion. Therefore, we propose several recommendations for systems that integrate conversational and web search.
@@
For tasks with closer psychological distances, we recommend returning a list of web search results. For example, when searching for local restaurants or immediate weather forecasts, web search can provide timely and specific information that directly meets users’ needs. However, for tasks with a greater psychological distance, we suggest providing synthesized information generated by conversational search. For example, when exploring potential travel destinations or learning about new technologies, conversational search can provide a comprehensive information framework to help users build a preliminary understanding of the topic.
Our findings also suggest that credibility perceptions, influenced by psychological distance, can guide how combined IR systems present and contextualize information. For tasks with closer psychological distances, where users may prioritize source authenticity, systems should emphasize source attributions in web search results, enhancing perceived credibility through transparency and trustworthiness. Conversely, for tasks with greater psychological distances, where users value abstract and synthesized information, conversational search could focus on providing high-level overviews while clearly communicating the limitations of LLM-generated content, such as potential biases or incomplete source references.
@@
Although past HCI research did not focus on the psychological distance of tasks, it explored factors influencing user choices between conversational and web search in specific task scenarios. A study on programming tasks noted that programmers opt for conversational search when objectives are clear, seeking results aligned with their goals (Yen et al., 2024). This phenomenon, stemming from the confirmatory nature of conversational search, known as the echo chamber effect, has been extensively studied in HCI (Sharma et al., 2024; Tanprasert et al., 2024).
To optimize this, some studies have suggested designs that allow conversational systems to actively present different viewpoints (Feltwell et al., 2020) and designs that proactively remind users of current information biases in conversations (Sharma et al., 2024). However, we propose that integrating conversational and web search, guided by psychological distance, offers a more holistic solution. By dynamically adapting the balance between the two modes based on user needs, systems can mitigate biases while enhancing the diversity and relevance of presented information. Additionally, HCI research discussing user trust in ChatGPT versus web search noted that generative AI’s lack of information sources reduces user trust (Jung et al., 2024). Yet, its assistance in ideation encourages its use. Our participants expressed similar views  [^11] Specific data are presented in the Appendix (A.3), uploaded with Supplementary Material., preferring to obtain preliminary information through conversational search, then verify this information and acquire more details through web search.
@@ 6.4. Limitations
This study has several limitations. First, our study only considered information seeking tasks in travel scenarios. These tasks, being general in nature, may have influenced user perceptions of LLM-powered conversational search, which is better suited for general rather than specialized tasks. This limitation might obscure the specific differences in user perceptions between web search and LLM-powered conversational search. For instance, we observed no significant differences in most metrics when the psychological distance was near. This pattern may not apply to other scenarios and also prevented us from further exploring the specific impacts of tasks with near psychological distance on both types of IR tools. However, with web search as a control, we found no significant differences in most ratings between the two IR tools when the psychological distance was near, but significant differences emerged when the psychological distance was far. This outcome still addresses our main research question and reveals a general principle: Users prefer LLM-powered conversational search when the psychological distance is greater.
Furthermore, our study did not counterbalance the order of conditions, as web search was always presented first, followed by conversational search. This decision was made for two primary reasons. First, conversational search may lead participants to use a wider variety of prompts, resulting in significant variability in the results they obtain. If conversational search were presented first, the knowledge gained during this task could significantly influence the behavior of participants in the subsequent task. In contrast, web search is a more familiar tool, and participants are more likely to exhibit consistent behaviors and obtain relatively similar results during this task. By placing web search first, we aimed to create a pseudo-control condition, ensuring greater consistency across participants when they transitioned to the conversational search condition. Second, any potential bias introduced by the fixed order of conditions would therefore be consistent across all experimental groups, which may not compromise the validity of our primary findings. That said, the fixed order remains a limitation, as it may have influenced the effectiveness of our psychological distance manipulations. For instance, participants might have adjusted their perceptions of psychological distance after completing the first task (web search), which could have affected their subsequent interactions with conversational search. Future studies could adopt a counterbalanced design to explore how task order interacts with psychological distance manipulations and influences user perceptions and preferences.
Finally, the custom search interface we implemented does not fully represent a real, commercial search page; it lacks features like image previews and related searches, especially compared to the actual Google Search website. However, since the search results are sourced from the genuine Google Search API, we consider them to still fairly represent a real search interaction.